OpenStudy (anonymous):

Suppose csc x = 2, cos x > 0 find tan, cot, sec, csc, sin, cos

myininaya (myininaya):

why is this not working :(

OpenStudy (anonymous):

using the unit circle.

myininaya (myininaya):

yes its not showing everything i posted :(

OpenStudy (turingtest):

1/sinx=2 sinx=1/2 x=pi/6 plug in pi/6 to each function and take the positive value because cosx>0

myininaya (myininaya):

\[\csc(x)=\frac{hyp}{opp}=\frac{2}{1}\] \[=> \sin(x)>0 \] since \[\sin(x)=\frac{1}{2}\]  so we are working in quadrant 1

myininaya (myininaya):

what are you guys seeing in my post?

myininaya (myininaya):

all i see is so we are working in quadrant 1

OpenStudy (anonymous):

yes thats all you're posting lol

OpenStudy (turingtest):

all I see is the same

myininaya (myininaya):

:( let me try this again

OpenStudy (turingtest):

so we are working in quadrant 1 you say?

myininaya (myininaya):

\[\csc(x)=2 => \sin(x)=\frac{1}{2} >0\] we want cos(x) and sin(x) to both be positive  \[\cos(x)=\frac{\sqrt{3}}{2}\]